# Home Swapping – An Innovative Approach to Reduce Traffic Congestion and Carbon Emissions

### v1

@@ Abstract
Urban traffic congestion, worsened by the rapid urbanization and the increasing prevalence of private vehicles, has significantly increased commuting time for everyone. In this paper, we used a dataset with over 400,000 real mobility trajectories of individuals spanning 9 days in a major Chinese city to investigate an innovative approach to swap homes between households in addressing the challenge of peak-hour traffic congestion. We observed that, empirically, households choose their home location strategically such that the average commuting distance is roughly $3$ times less than that when their home is randomly located, showing features of self-organization. Remarkably, we found that the average commuting distance can be further reduced by $50\%$ through home swapping at the city-level, leading to a large reduction in traffic congestion. To make home swapping more realistic, we swap homes only if the following socio-demographic factors including the distance from the city center, housing price and amenity accessibility are preserved for both households, such that the average commuting distance can still be reduced by $13\%$. As both home-workplace distance and traffic congestion are reduced, as a side benefit, carbon emissions from vehicles are also greatly reduced by almost $80\%$, and $40\%$ when socio-demographic factors are considered. The distance from the city center is shown to be the most influential factor affecting the benefit brought by home swapping, and further analysis indicates that developing a polycentric city layout could significantly enhance such benefit. This study suggests that mitigating traffic congestion requires a long-term, holistic and strategic approach to urban planning, suggesting a need for coordinating individual residence locations and a polycentric city layout.
@@ I Introduction
With the rapid urbanization, an increasing ratio of the world population resides in urban areas sun2020dramatic . While the room for expansion in urban road infrastructure is small to cope with the increasing population, traffic congestion becomes more severe ccolak2016understanding ; pi2019visual ; zhou2022identifying ; dasgupta2021spatiotemporal ; wang2023aggravated ; manfredi2018mobility ; dahlmanns2023optimizing ; zhang2019scale . This does not only reduce the daily comfort of city commuters, but also waste a lot of resources in terms of time and energy, and induce air and noise pollution, impacting the socio-economy of cities at a holistic level.
@@
As we all know, the root cause of traffic congestion during peak hours lies in the need of individuals to travel between home and their workplace or school  liu2018investigating . The longer the commuting distance, the more likely private motor vehicles are used, which leads to more vehicles on roads, and hence increased traffic congestion during peak hours  wheaton2004commuting ; zhao2020long . On the contrary, if the commuting distance is short, people can use non-motor vehicles such as bicycles or even walk to their workplace. Thus, there were studies which examine residences, workplaces, and commuting distances using empirical datasets  huang2018tracking ; plaut2006intra ; gerber2020workplace ; surprenant2013commuting ; guo2021fractal ; suzuki2012jobs ; zhao2011impact ; asmussen2024interplay ; hu2017different ; ta2017understanding ; shen2021job . Studies show that moving or changing jobs can change the spatial arrangement of residences and workplaces for single-commuter and dual-commuter households, reducing their commuting distance and time  huang2018tracking ; surprenant2013commuting . However, these studies were usually based on surveys, which have a finite number of respondents and hence do not fully reveal the benefit brought by reallocation of residence at the city level.
@@
In this paper, we examine the impact of city-wide home-swapping in reducing traffic congestion and alleviating carbon emission. We used more than 400,000 real individual mobility trajectories over 9 days from a Chinese city, to infer the home and workplace location as well as household members of each individual, and identify all the commuter path during peak hours using Baidu map BaiduPlatform . We found that randomly re-shuffling residence locations increases the average commuting distance by $2.83$ times, indicating a strong self-organization in the actual choices of residence location. By strategic home swapping, urban congestion during peak hours is greatly reduced since the average commuting distance is reduced by $50.4\%$, which complies with the notion of a 15-minute city abbiasov202415 ; khavarian202315 ; papadopoulos2023measuring . Our study is the first to verify the benefit brought by home-swapping using such a real and big data.
@@
To make home swapping more realistic, we swap home only if the following three socio-demographic factors are preserved for both households, namely the distance from the city center, housing price and amenity accessibility. In this case, the average commuting distance can still be reduced by 13%. Owing to the combined effect of shorter commuting distance and reduced traffic congestion, carbon emissions from vehicles are also greatly reduced by $78\%$, and by $43\%$ when socio-demographic factors are considered. Among these, the distance from the city center influences home-swapping most, followed by the amenity accessibility and housing price. Based on our findings, we found that establishing a polycentric, i.e. multi-center, city layout can effectively increase the benefit brought by home swapping.
@@ II Results
Our big data of empirical human trajectory are based on the 4G communication records between base stations and mobile phones provided by one of the three major service providers in Shijiazhuang, the capital of Hebei province in China, from May $22^{\rm nd}$ to $28^{\rm th}$, 2017, and from June $3^{\rm rd}$ to $4^{\rm th}$. The original data include records from around 3 million users out of a total population of 10 million in the city and spans 9 days, including 2 weekends and 5 weekdays. To identify meaningful locations in the user trajectories, we divide the whole period covered by the dataset into 15-minute intervals (see Materials and Methods). If an individual stayed in the same location for more than 10 minutes in each interval, the location was recorded as a stop for that user. In addition, when we compute the similarity between the trajectories of individuals later, the division into time intervals facilitates the comparison of defined stops of individuals in each time interval. Other than this high-frequency mobility dataset in Shijiazhuang, we also used the dataset of residence and workplace from Shenzhen of China and the dataset of call detail records (CDR) from Abidjan of Cote d’Ivoire to cross-validate the results observed in Shijiazhuang (see Section S1 of the Supplementary Information (SI) for details).
@@
Next, we infer the household membership between individuals. Since (1) members in the same household usually have different workplace locations but they may enjoy the weekend outing together, and (2) previous studies demonstrated that the intimacy between individuals is related to the similarity between their mobility trajectories  wang2011human , we use the similarity between the trajectories of two individuals with the same residence location from the 4 days in the 2 weekends to evaluate their household membership. Specifically, we denote the similarity between the mobility trajectory of individual $i$ and $j$ in a single day by $S_{ij}$ (see Materials and Methods). The closer the value of $S_{ij}$ to $1$, the more similar their trajectory, and closer the relationship between $i$ and $j$. Nevertheless, even in weekends, family members may only spend part of the day together rather than the entire day. Therefore, we determine if $i$ and $j$ are household members by using the highest $S_{ij}$ among the four values found in each of the 4 days of the 2 weekends. If $S_{ij}>0.5$, individual $i$ and $j$ are considered members in the same household. A schematic example of computing the similarity between the trajectories of two individuals is shown in Fig. 1A.
@@
To start analyzing the potential benefit brought by home swapping, we first show the spatial distribution of the location of residence and workplace of commuters in the city in Fig. 2A and Fig. 2B respectively. We divide the urban area of Shijiazhuang into a grid of squares with a dimension of $500m\times 500m$; the more the residence or workplace locations in a grid, the higher the intensity of the color on the heat maps in Figs. 2A and 2B. As we can see, areas close to the city center (e.g. Yuhua ,Chang’an ,Xinhua and Qiaoxi Districts) have a high concentration of residence and workplace locations, and the opposite is observed in the surrounding suburban areas. Residence locations are more dispersed and numerous, while workplaces are more concentrated near the city center (see Fig. 2C). The KL divergence (see Materials and Methods) between the two spatial distributions is 0.096, indicating that they are very similar. Similar observations are obtained in Shenzhen and Abidjan and are shown in Fig. S16 and S20 respectively.
Despite the similarity between the overall spatial distributions of residence and workplace locations, it does not necessarily imply that commuters live near their workplaces. As shown in Fig. 2D, the distribution of their daily commuting distance is roughly exponential, such that many commuters commute a long distance daily, suggesting a significant room for exploitation. To examine whether commuters used motor vehicles for commuting, we found that the travel distance of Shijiazhuang residents on-demand vehicles peaked at 2.5 km (see in Fig. S1). We therefore assume that 2.5 km is the shortest distance for commuters to use motor vehicles for commuting. In addition, we also shown in Sec S6 of the SI a different threshold does not qualitatively affect the obtained results. Based on the empirical home-workplace distance, almost $64\%$ of the Shijiazhuang citizens commute a distance larger than 2.5 km and thus they are assumed to use motor vehicles for commuting. In addition, we see that the inferred departure time of commuters from their home peaks within a short period during the morning rush hours in Fig. S2, and together with the long commuting distance, are the dominant factor for urban traffic congestion according to  wheaton2004commuting ; yildirimoglu2021staggered .
@@
Interestingly, we note a second peak at an intermediate $d_{\rm max}-d_{\rm min}$, implying that many other households reside at a location closer to one of the household members. As shown in Fig. 2F, dual-commuter households exhibit a clear bifurcation, i.e. some households choose to live in between workplaces such that $(d_{\rm max},d_{\rm min})$’s lie near the diagonal, while others opt for locations close to one member but farther away from the other with $(d_{\rm max},d_{\rm min})$’s lying near the horizontal axis. With the increase in the number of household members, it becomes less feasible for families to live close to everyone’s workplaces. In this case, we see that $(d_{\rm max},d_{\rm min})$’s are mainly found near the horizontal axis, implying that they are more inclined to reside in locations close to the workplaces of a few members, sacrificing at least one member’s commuting time, consistent with the conclusion in  surprenant2013commuting .
@@
Greedy Home Swapping (GHS). To examine the benefit brought by strategic home swapping, we propose a greedy strategy to swap homes only if the total commuting distance of the two swapped households decreases. Due to the restriction in living space, we only allow home-swapping between 1-commuter and 2-commuter households, as well as swapping between 3-commuter and 4-commuter households. We terminate the swapping attempts if no households can be swapped to reduce commuting distance for a consecutive $1\times 10^{6}$ attempts. As the goal of home-swapping is effectively to match each household to a residence location given the workplace locations of all citizens, the Hungarian algorithm is able to identify the optimal solution for this matching problem and hence the optimal allocation of residence locations for all household. Nevertheless, its computation complexity of $O(N^{3})$ and the large number of $2.4\times 10^{5}$ households in our real data would make the matching intractable. Therefore, we employed the above swapping procedure to reduce the commuting distance of the swapped households and the city as whole.
In this case, after home swapping, the decrease in commuting distance fundamentally reduces traffic congestion, through the combined effect of a shorter traveling distance and a reduced number of motor vehicles used since some commuters may walk or cycle instead. An example of the impact of greedy home swapping (GHS) on individual households is illustrated in Fig. 1C-D. In Fig. 3A, we further show the average commuting distance over all commuters in the city based on the real data, after random and greedy home swapping, which we denote by $\bar{d}_{\rm real}$, $\bar{d}_{\rm random}$ and $\bar{d}_{\rm greedy}$ respectively. As we discussed before, $\bar{d}_{\rm random}=2.83\bar{d}_{\rm real}$, and after greedy home swapping, $\bar{d}_{\rm greedy}=0.5\bar{d}_{\rm real}$. To better compare the benefits brought by different home swapping approaches later, we denote $\Delta d$ as the percentage decrease in the average commuting distance over all commuters, given by
@@
In Fig. 3B, we further show the scatter plots of the average commuting distances in a household before and after GHS. We find that the average commuting distance for most households are shortened. Nevertheless, some households with already a short commuting distance have their commuting distance increased after GHS, but remains within an acceptable range. It is because the households which swap home with these families have a larger reduction in commuting distances compared to their increases. This is also an example of individual optimum versus social optimum, such that some individual households have to sacrifice to choose a sub-optimal home location to benefit others. Finally, there are also a few households whose commuting distances have not changed, either because they did not swap home or the commuting distance between their new home and workplaces remain unchanged.
@@
After households’ residences are re-allocated, there are changes in their routes, potentially their modes of transportation, and departure times (see Materials and Methods). For instance, commuters with a commuting distance of less than 2.5 km may switch to a non-motorized travel such as bicycles. In this case, the congestion on different road segments at any specific time of the day may change.
@@
To quantify the change in congestion, we define a city-level congestion index, called the average congestion coefficient $I_{t}$ (see Materials and Methods). We then analyze this index at different times of the day based on the real dataset, and after random and greedy home swapping. In Fig. 3C, for all periods of time, we found that the average congestion coefficient after RHS is highest, while that after GHS is lowest, with the empirical one in between. Further analysis was conducted on the traffic condition during the morning peak hours of 7-8 AM, when most commuters leave home for work. In the case of RHS, most road segments are severely congested and vehicles move slowly. After GHS, only a few road segments remain congested as shown in Fig. 3D-F; based on these results, one can anticipate that GHS alleviate congestion throughout the rush hour (see Fig. S4).
To verify the above remarkable benefit brought by GHS in cities other than Shijiazhuang, we conducted similar simulations using datasets obtained from Shenzhen and Abidjan and obtained qualitatively similar results and benefits as in Fig. 3A (see SI section 7 for details). These results show that there are great benefits brought by GHS in the cities examined, which can be generic among all other cities.
@@
Greedy Home Swapping with Socio-demographic Consideration (SDGHS). While Greedy Home Swapping is based solely on reducing commuting distance which significantly alleviates traffic congestion, previous literature has confirmed that socio-demographic factors play a vital role in people’s choice of residence  tillema2010influence ; guo2020modeling ; rouwendal2001preferences ; ho2016workplace ; hu2017different ; guan2020multiplicity ; liu2022quantifying ; wang2024does ; al2024pathway ; cao2017examining ; blumenberg2023housing . Families, although not fond of spending time in transit, have preference for specific housing characteristics that are strong enough for them to accept a longer commuting time. In view of this socio-demographic consideration, to make home swapping more realistic, we further consider three factors during home swapping: (1) distance from the city center, (2) housing price, and (3) amenity accessibility.
@@
Specifically, on top of reducing commuting distance, two households swap their home only if
@@
    the difference in the distance between their home and the city center is less than $T_{\rm center}$ km (c-GHS) - It is because the distance from the city center is an important factor when families choose their home. According to ref. xu2023urban and the data, we define the most visited location in Shijiazhuang to be its city center, i.e. the largest central business distract , i.e. the Beiguo Shopping Mall in Shijiazhuang.
@@
    the difference between the housing prices of their home is less than $T_{\rm price}\%$ (p-GHS) - Since housing prices affect their choice of residence, for instance, they generally choose residence within their financial capacity. Here, we use the housing prices of those second-hand properties to represent housing price in that region (see Materials and Methods).
@@
    the difference between the availability of amenities around the two residences is less than $T_{\rm amenity}$ (a-GHS) - Since households consider the surrounding living environment when they choose their home abbiasov202415 , here we use the information of the Points of Interest (POI) around the corresponding residence to represent its living environment. For locations without the information of their surrounding POI (e.g. mostly villages), we do not swap their home with others (see Materials and Methods).
@@
In Fig. 4D, we further examine $\Delta d$ under the joint consideration of all the three socio-demographic factors with different level of tolerance, which we denote by Socio-demographic Greedy Home Swapping (SDGHS). As we can see, $\Delta d$ reaches $25\%$ under high tolerance, i.e. $T_{\rm center}=$ 1000 meters, $T_{\rm price}=25\%$ and $T_{\rm amenity}=0.5$, but achieves only $3\%$ under low tolerance, i.e. $T_{\rm center}=$ 100 meters, $T_{\rm price}=5\%$ and $T_{\rm amenity}=0.1$. In Fig. 4E, we compare the average congestion coefficient $I_{t}$ computed based on the inferred residence and workplace location from the real dataset and the path of each commuter recommended in Baidu Map during morning and evening rush hour, with those after home swapping under the forementioned socio-demographic consideration with different tolerances. We found that the status of $I_{t}$ under strict consideration (low tolerance) of the factors is closer to that in the real data, while the congestion under loose consideration (high tolerance) has improved as shown in Fig. 4E. $I_{t}$ under medium consideration (medium tolerance) in rush hours is shown in Fig. S5. We also conducted similar tests using datasets obtained from Shenzhen and Abidjan, and obtained qualitatively similar results (see Fig. S18, S19A and S23A).
@@
Finally, in addition to the average congestion coefficient, we analyze the degree of congestion on each road after GHS taking different factors into account, to provide a clearer understanding on how traffic congestion is relieved. As shown in Fig. S6, home swapping can effectively reduce the number of congested and slow-moving road segments during the rush hour. In addition, we show in Fig. 4F that there is a larger reduction of road congestion between the empirical congestion and that after home swapping under the consideration of housing price and amenity accessibility with medium tolerance. However, after GHS taking into account the distance to the city center, some roads remain similarly congested, implying that the distance to the city center is the main consideration of households which hinder the reduction of city-wide commuting distance by GHS. Further comparison among the reductions in average commuting distance based on any 2 out of the 3 socio-demographic factors is shown in Fig. S7, and the results also suggest $T_{\rm center}$ emerges as the major factor influencing GHS to improving urban congestion.

### v2

# Guiding Self-Organizing Dynamics of Residential Choice in Cities to Reduce Traffic Congestion and Carbon Emissions

@@ Abstract
Rapid urbanization and growing vehicle ownership exacerbate traffic congestion and prolong commute times. We examine the self-organizing dynamics of residential choice via a hypothetical home-swapping process to mitigate peak-hour traffic congestion and carbon emissions. Specifically, we analyze over $400,000$ trajectories from $9$ days in a major Chinese city, revealing that actual average commuting distance is approximately three times shorter than under random residential distribution, indicating significant self-organization. Notably, city-wide home swapping reduces commuting distance by $50.4\%$, substantially easing traffic congestion, thereby reducing carbon emissions by $77.3\%$. Even with the consideration of socio-demographic factors and individual needs, the reductions remain significant: $8.1\%$-$10.3\%$ in commuting distance and $27.4\%$-$34.4\%$ in carbon emissions. Considering the potential induction of additional non-commuting trips, the reduction in carbon emissions remains substantial. Given the primacy of distance to the city center, polycentric city layouts can enhance these benefits. For validation, we use another dataset covering China’s 28 major cities to confirm these findings. Finally, we introduce a data-driven model to elucidate self-organizing dynamics of residential choice and analyze the feasibility of government coordination. These insights demonstrate that a synergistic alignment of residential choices can leverage individual and city-level benefits, effectively alleviating commuting congestion and associated emissions.
@@ I Introduction
With the rapid urbanization, an increasing ratio of the world population resides in urban areas sun2020dramatic . While the room for expansion in urban road infrastructure is small to cope with the increasing population, traffic congestion becomes more severe ccolak2016understanding ; pi2019visual ; zhou2022identifying ; dasgupta2021spatiotemporal ; wang2023aggravated ; manfredi2018mobility ; dahlmanns2023optimizing ; zhang2019scale . This does not only reduce the daily comfort of city commuters, but also wastes a lot of resources in terms of time and energy, and induce air and noise pollution, impacting the socio-economy of cities at a holistic level.
@@
As we all know, the root cause of traffic congestion during peak hours lies in the need of individuals to travel between home and their workplace or school liu2018investigating ; NIU2025105583 ; 2017transportreport . The longer the commuting distance, the more likely private motor vehicles are used, which leads to more vehicles on roads, and hence increased traffic congestion during peak hours wheaton2004commuting ; zhao2020long . On the contrary, if the commuting distance is short, people can use non-motor vehicles such as bicycles or even walk to their workplace. Thus, there were studies which examine residences, workplaces, and commuting distances using empirical datasets huang2018tracking ; plaut2006intra ; gerber2020workplace ; surprenant2013commuting ; guo2021fractal ; suzuki2012jobs ; zhao2011impact ; asmussen2024interplay ; hu2017different ; ta2017understanding ; shen2021job . Studies show that moving or changing jobs can change the spatial arrangement of residences and workplaces for single-commuter and dual-commuter households, reducing their commuting distance and time huang2018tracking ; surprenant2013commuting . Although well-designed surveys can indeed provide statistically valid insights into residential preference through carefully constructed samples yan2013diversity , most existing studies were usually based on surveys from a small number of respondents, and to our knowledge, have never revealed the benefit brought by reallocation of residence at the city level.
@@
In this paper, we examine the impact of a hypothetical city-wide home-swapping in reducing traffic congestion and alleviating carbon emissions. We used more than 400,000 real individual mobility trajectories over 9 days from a Chinese city, Shijiazhuang, to infer the home and workplace location as well as household members of each individual, and identify all the commuter path during peak hours using Baidu map BaiduPlatform . We found that randomly re-shuffling residence locations increases the average commuting distance by $2.83$ times, indicating a strong self-organization in the actual choice of residence location. The self-organization bak1988self ; haken2006information in our context refers to the emergent spatial patterns of residential locations that arise from decentralized household decisions made based on their local information, without centralized coordination, such that the average commuting distance is still much smaller than that of the random residential distribution. By strategic home swapping, urban congestion during peak hours is greatly reduced since the average commuting distance is reduced by $50.4\%$, which complies with the notion of a 15-minute city abbiasov202415 ; khavarian202315 ; papadopoulos2023measuring .
@@
To make home swapping more realistic, we consider the heterogeneity of households and swap homes only if the three socio-demographic factors including the distance to the city center, housing price and amenity accessibility, and needs of individual household members are preserved for both households. In this case, the average commuting distance can be reduced by $13.4\%$ if only socio-demographic factors are considered. By also considering the needs of individual household members, the average commuting distance can still be reduced by $8.1\%$ to $10.3\%$. In all these cases, owing to the combined effect of shorter commuting distance and reduced traffic congestion, carbon emissions from vehicles are also greatly reduced by $77.3\%$ without constraints and by $27.4\%$ to $34.4\%$ when socio-demographic factors and the needs of individual household members are considered, respectively. Among the three socio-demographic factors, the distance to the city center influences home-swapping most, followed by the amenity accessibility and housing price. Based on our findings, we found that establishing a polycentric, i.e. multi-center, city layout can effectively increase the benefit brought by guiding self-organizing residential choice.
While a great benefit of home-swapping has been shown in the simulation on the dataset of Shijiazhuang, we further verify the universality of our results in cities other than Shijiazhuang and investigate the impact of induced traffic by the reduced traffic congestion. Specifically, we examine the impact of home swapping using another origin-destination dataset covering 28 major cities in China zhao2024unravelling , and the results are consistent with our findings in Shijiazhuang, suggesting universal benefits can be brought by home swapping. Since traffic congestion is extensively reduced by home swapping, non-commuting traffic such as leisure trips may be induced. Drawing on trajectory data from Shijiazhuang, randomly adding traffic during 8-10 a.m. to the traffic at peak hour (7-8 a.m.) after home swapping, we found that the reduction in carbon emissions remains substantial. At last, we introduce a data-driven model to explore the self-organizing dynamics of residential choice based on urban economics theory. By introducing government coordination to the model, we can obtain a similar result in the reduction in average commuting distance as in GHS and SDGHS, which suggests a feasible approach to achieve a similar residential allocation as in home-swapping.
@@ II Results
Our big data of empirical human trajectory are based on the 4G communication records between base stations and mobile phones provided by one of the three major service providers in Shijiazhuang, the capital of Hebei province in China, from May $22^{\rm nd}$ to $28^{\rm th}$, 2017, and from June $3^{\rm rd}$ to $4^{\rm th}$. The original data include records from around 3 million users out of a total population of 10 million in the city and span 9 days, including 2 weekends and 5 weekdays. To identify meaningful locations in the user trajectories, we divide the whole period covered by the dataset into 15-minute intervals (see Materials and Methods). If an individual stayed in the same location for more than 10 minutes in each interval, the location was recorded as a stop for that user. In addition, when we compute the similarity between the trajectories of individuals later, the division into time intervals facilitates the comparison of defined stops of individuals in each time interval. Other than this high-resolution mobility dataset in Shijiazhuang, we also studied the origin-destination dataset at peak hour (i.e. 7-8 a.m.) of 28 major cities in China to verify the results observed in Shijiazhuang (see Section S1 of the Supplementary Information (SI) for details).
@@
Next, we infer the household membership between individuals. Since (1) members in the same household usually have different workplace locations but they may enjoy the weekend outing together, and (2) previous studies demonstrated that the intimacy between individuals is related to the similarity between their mobility trajectories wang2011human , we use the similarity between the trajectories of two individuals with the same residence location from the 4 days in the 2 weekends to evaluate their household membership. Specifically, we denote the similarity between the mobility trajectory of individual $i$ and $j$ in a single day by $S_{ij}$ (see Materials and Methods). The closer the value of $S_{ij}$ to $1$, the more similar their trajectories, and closer the relationship between $i$ and $j$. Nevertheless, even in weekends, family members may only spend part of the day together rather than the entire day. Therefore, we determine if $i$ and $j$ are household members by using the highest $S_{ij}$ among the four values found in each of the 4 days of the 2 weekends. If $S_{ij}\geq 0.5$, individual $i$ and $j$ are considered members in the same household. A schematic example of computing the similarity between the trajectories of two individuals is shown in Fig. 1A.
@@
To start analyzing the potential benefit brought by home swapping, we first show the spatial distribution of the location of residence and workplace of commuters in the city in Fig. 2A and Fig. 2B respectively. We divide the urban area of Shijiazhuang into a grid of squares with a dimension of $500m\times 500m$; the more the residence or workplace locations in a grid, the higher the intensity of the color on the heat maps in Figs. 2A and 2B. As we can see, areas close to the city center (e.g. Yuhua ,Chang’an ,Xinhua and Qiaoxi Districts) have a high concentration of residence and workplace locations, and the opposite is observed in the surrounding suburban areas. Residence locations are more dispersed and numerous, while workplaces are more concentrated near the city center (see Fig. 2C). The KL divergence (see Materials and Methods) between the two spatial distributions is 0.096, indicating that they are very similar.
Despite the similarity between the overall spatial distributions of residence and workplace locations, it does not necessarily imply that commuters live near their workplaces. As shown in Fig. 2D, the distribution of their daily commuting distance is roughly exponential, such that many commuters commute a long distance daily, suggesting a significant room for exploitation. To examine whether commuters used motor vehicles for commuting, we found that the travel distance of Shijiazhuang residents on-demand vehicles peaked at 2.5 km (see in Fig. S1). We therefore assume that 2.5 km is the shortest distance for commuters to use motor vehicles for commuting. Based on the empirical home-workplace distance, almost $64\%$ of the Shijiazhuang citizens commute a distance larger than 2.5 km and thus they are assumed to use motor vehicles for commuting. In other words, commuters switch from motorized travel to non-motorized travel encompassing walking, cycling, and e-bikes when the commuting distance is less than 2.5 km. This switching threshold of 2.5 km is in line with classification of non-motorized and motorized travel in the 2020 National Commuting Report of China 2020tongqinbaogao . In addition, we also shown in Sec. S7 of the SI that different thresholds do not qualitatively affect the obtained results.
In addition, we see that the inferred departure time of commuters from their home peaks within a short period during the morning rush hours in Fig. S2, and together with the long commuting distance, are the dominant factor for urban traffic congestion according to ref. wheaton2004commuting ; yildirimoglu2021staggered .
@@
Interestingly, we note a second peak at an intermediate $d_{\rm max}-d_{\rm min}$, implying that many other households reside at a location closer to one of the household members. As shown in Fig. 2F, dual-commuter households exhibit a clear bifurcation, i.e. some households choose to live in between workplaces such that $(d_{\rm max},d_{\rm min})$’s lie near the diagonal, while others opt for locations close to one member but farther away from the other with $(d_{\rm max},d_{\rm min})$’s lying near the horizontal axis. With the increase in the number of household members, it becomes less feasible for families to live close to everyone’s workplaces. In this case, we see that $(d_{\rm max},d_{\rm min})$’s are mainly found near the horizontal axis, implying that they are more inclined to reside in locations close to the workplaces of a few members, sacrificing at least one member’s commuting time, consistent with the conclusion in ref. surprenant2013commuting .
@@
Greedy Home Swapping (GHS). To examine the benefit brought by strategic home swapping, we propose a greedy strategy to swap homes only if the total commuting distance of the two swapped households decreases. Due to the restriction in living space, we only allow home-swapping between 1-commuter and 2-commuter households, as well as swapping between 3-commuter and 4-commuter households. We terminate the swapping attempts if no households can be swapped to reduce commuting distance for a consecutive $1\times 10^{6}$ attempts. As the goal of home-swapping is effectively to match each household to a residence location given the workplace locations of all citizens, the Hungarian algorithm is able to identify the optimal solution for this matching problem and hence the optimal allocation of residence locations for all households. Nevertheless, its computation complexity of $O(N^{3})$ and the large number of $2.4\times 10^{5}$ households in our real data would make the matching intractable. Therefore, we employed the above swapping procedure to reduce the commuting distance of the swapped households and the city as whole.
In this case, after home swapping, the decrease in commuting distance fundamentally reduces traffic congestion, through the combined effect of a shorter traveling distance, more dispersed departure times, and a reduced number of motor vehicles used since some commuters may walk or cycle instead. An example of the impact of Greedy Home Swapping (GHS) on individual households is illustrated in Fig. 1C-D. In Fig. 3A, we further show the average commuting distance over all commuters in the city based on the real data, after random and greedy home swapping, which we denote by $\bar{d}_{\rm real}$, $\bar{d}_{\rm random}$ and $\bar{d}_{\rm greedy}$ respectively. As we discussed before, $\bar{d}_{\rm random}=2.83\bar{d}_{\rm real}$, and after GHS, $\bar{d}_{\rm greedy}=0.5\bar{d}_{\rm real}$. To better compare the benefits brought by different home swapping approaches later, we denote $\Delta d$ as the percentage decrease in the average commuting distance over all commuters, given by
@@
In Fig. 3B, we further show the scatter plots of the average commuting distances in a household ($\bar{d}_{\rm household}$) before and after GHS. We find that the average commuting distance for most households are shortened. Based on our empirical data, the mean of $\bar{d}_{\rm household}$ is 8.3 km. After GHS, only $6.1\%$ of households have their average commuting distance increased, and just $0.34\%$ of all households have their average commuting distance increased by more than 5 km. It is because the households which swap homes with these families have a larger reduction in commuting distances compared to their increases. This is also an example of individual optimum versus social optimum, such that some individual households have to sacrifice to choose a sub-optimal home location to benefit others. Finally, there are also a few households whose commuting distances have not changed, either because they did not swap homes or the commuting distance between their new home and workplaces remain unchanged.
@@
After households’ residences are re-allocated, there are changes in their routes, potentially their modes of transportation, and departure time (see Materials and Methods). For instance, commuters with a commuting distance of less than 2.5 km may switch to a non-motorized travel such as bicycles. In this case, the congestion on different road segments at any specific time of the day may change.
@@
To quantify the change in congestion, we define a city-level congestion index, called the average congestion coefficient $I_{t}$ (see Materials and Methods). We then analyze this index at different time of the day based on the real dataset, and after random and greedy home swapping. In Fig. 3C, for all periods of time, we found that the average congestion coefficient after RHS is highest, while that after GHS is lowest, with the empirical one in between. Further analysis was conducted on the traffic condition during the morning peak hour of 7-8 a.m., when most commuters leave home for work. In the case of RHS, most road segments are severely congested and vehicles move slowly. After GHS, only a few road segments remain congested as shown in Figures 3D-F; based on these results, one can anticipate that GHS alleviate congestion throughout the rush hours (see Fig. S4).
@@
Greedy Home Swapping with Socio-demographic Consideration (SDGHS). While Greedy Home Swapping is based solely on reducing commuting distance which significantly alleviates traffic congestion, previous literature has confirmed that socio-demographic factors play a vital role in people’s choice of residence tillema2010influence ; guo2020modeling ; rouwendal2001preferences ; ho2016workplace ; hu2017different ; guan2020multiplicity ; liu2022quantifying ; wang2024does ; al2024pathway ; cao2017examining ; blumenberg2023housing .
Households, although not fond of spending time in transit, have preference for specific housing characteristics that are strong enough for them to accept a longer commuting time. To identify the difference of each family in capturing economic and locational preferences, we use (1) the distance to the city center, (2) housing price, and (3) amenity accessibility, to serve as proxies to characterize household-level heterogeneity. These metrics were selected because they reflect household-specific attributes rather than purely spatial characteristics, enabling home swapping to take into account the major consideration of individual household while keeping the self-organizing principles of the original GHS framework.
@@
Specifically, on top of reducing commuting distance, two households swap their homes only if
@@
    the difference in the distance between their home and the city center is less than $T_{\rm center}$ km (c-GHS) - c-GHS portrays the centrality consideration of residential choice. It is because the distance to the city center is an important factor when families choose their homes. According to ref. xu2023urban and the data, we define the most visited location in Shijiazhuang to be its city center, i.e. the largest central business district , i.e. the Beiguo Shopping Mall in Shijiazhuang.
@@
    the difference between the housing prices of their home is less than $T_{\rm price}(\%)$ (p-GHS) - p-GHS incorporates housing price to model the financial consideration of households. Since housing price reflects the economic capacity of the household and affect their choice of residence, for instance, they generally choose residence within their financial capacity. Here, we use the city-wide housing transaction data of second-hand properties to describe the average housing price in each region (see Materials and Methods).
@@
    the difference between the availability of amenities around the two residences is less than $T_{\rm amenity}$ (a-GHS) - a-GHS captures the neighborhood consideration via the Points of Interest (POI) density. Since households consider the surrounding living environment when they choose their homes abbiasov202415 , here we use the information of POI around the corresponding residence to represent its living environment. For locations without the information of their surrounding POI (e.g. mostly villages), we do not swap their homes with others (see Materials and Methods).
@@
In Fig. 4D, we further examine $\Delta d$ under the joint consideration of all the three socio-demographic factors with different level of tolerance, which we denote by Socio-demographic Greedy Home Swapping (SDGHS). As we can see, $\Delta d$ reaches $25\%$ under high tolerance, i.e. $T_{\rm center}=1km$, $T_{\rm price}=25\%$ and $T_{\rm amenity}=0.5$, but achieves only $3\%$ under low tolerance, i.e. $T_{\rm center}=0.1km$, $T_{\rm price}=5\%$ and $T_{\rm amenity}=0.1$. In Fig. 4E, we compare the average congestion coefficient $I_{t}$ computed based on the inferred residence and workplace location from the real dataset and the path of each commuter recommended in Baidu Map during morning and evening rush hours, with those after home swapping under the forementioned socio-demographic consideration with different tolerances. We found that the status of $I_{t}$ under strict consideration (low tolerance) of the factors is closer to that in the real data, while the congestion under loose consideration (high tolerance) has improved as shown in Fig. 4E. $I_{t}$ under medium consideration (medium tolerance), $T_{\rm center}=0.5km$, $T_{\rm price}=15\%$ and $T_{\rm amenity}=0.3$, in rush hours is shown in Fig. S5, and over $80\%$ of households benefit under medium tolerance conditions (Fig. S19). For the sake of simplicity in discussion about SDGHS, we used SDGHS with medium tolerance for simulation in our subsequent analysis. In addition to the average congestion coefficient, we analyze the degree of congestion on each road after GHS taking different factors into account, to provide a clearer understanding on how traffic congestion is relieved. As shown in Fig. S6, home swapping can effectively reduce the number of congested and slow-moving road segments during the rush hour.
@@
In Fig. 4F, it is shown that there is a larger reduction of road congestion between the empirical congestion and that after home swapping under the consideration of housing price and amenity accessibility. However, after GHS taking into account the distance to the city center, some roads remain similarly congested, implying that the distance to the city center is the main consideration of households which hinder the reduction of city-wide commuting distance by GHS. Further comparison among the reductions in average commuting distance based on any 2 out of the 3 socio-demographic factors is shown in Fig. S7, and the results also suggest $T_{\rm center}$ emerges as the major factor influencing GHS to improving urban congestion.